OpenStudy (blackops2_143):

what is (n-5) x 6 ------- n divided by 4 when N= 8

OpenStudy (blackops2_143):

so is it 18 over 2

OpenStudy (blackops2_143):

it can't be a fraction. please help

OpenStudy (123ab456c):

6nx−12x

563blackghost (563blackghost):

So does it look like this....\[\frac{ (n-5) \times 6 }{ n \div 4 }\]

OpenStudy (blackops2_143):

yes

563blackghost (563blackghost):

Great ^^ so we would input 8 where n is first....\[\frac{ (8-5) \times 6 }{ 8 \div 4 }\] So we first would work on the numerator....so \[(8-5) \times 6\] lets first work on the paranthesis so we do 8 -5...

OpenStudy (blackops2_143):

3

OpenStudy (blackops2_143):

then 3 x 6 right?

563blackghost (563blackghost):

Correct ^^ So we get 3....and Correct once again...so what would 3 by 6 equal?

OpenStudy (blackops2_143):

18

563blackghost (563blackghost):

Yup :) Now we would do the denominator next.... so 8 divided by 4...

OpenStudy (blackops2_143):

2

563blackghost (563blackghost):

Correct ^^ so we get the fraction \[\frac{ 18 }{ 2 }\]Now in order to get your whole number we would divide these two so \[18 \div 2 =?\]

OpenStudy (blackops2_143):

9?

563blackghost (563blackghost):

Correct ^^ So the answer is 9 xD Great Job!

OpenStudy (blackops2_143):

thank you so much!

563blackghost (563blackghost):

np xD if you ever need help in the future and see me on just tag me in your question ^^